Elementary Functions and Calculus III Math 133 (Sec 42), Spring 2005 Course Document

Lecturer: David Rule Office: Eckhart 16 Phone: (773) 702 6045 Email: rule@math.uchicago.edu Website: www.math.uchicago.edu/˜ rule/133/

Lectures: Monday, Wednesday & Friday, 11:30am-12:20pm

Tutorials: Tuesday & Thursday, 12:00-1:20pm Marina Danilevsky (C 207) Daniel Clinton (GB 506) Jason Lau (GB 502)

Office hours: Wednesday & Thursday from 1:30pm to 2:30pm and by appointment.

Text: Calculus, Varberg, Purcell & Rigdon, Chapters 7, 8, 9 & 10.

Syllabus: The 131-2-3 sequence aims to teach the basic concepts and techniques of rigorous mathematics and apply these to certain problems. This course will cover basic differential equations, integration by parts, trigonometric substitution, l’Hôpitals rule, the improper Riemann integral, power series and Taylor series.

Assessment: Your grade will be determined by two in-class midterm tests, a final, and homeworks. I may occasionally give quizzes but these will not count towards your grade. The final constitutes 40% of your grade, the midterms 20% each, and homeworks total 20%.

The Final Examination: The final examination is on Monday, 6th June from 10:30am to 12:30pm. It is the policy of the Department of Mathematics that the following rules apply to final exams in all undergraduate mathematics courses:

1. The final exam must occur at the time and place designated on the College Final Exam Schedule. In particular, no final examinations may be given during the tenth week of the quarter, except in the case of graduating seniors.

2. Any student who wishes to depart from the scheduled final exam time for the course must receive permission from Paul Sally (office is Ry 350, phone is 2-7388, email is sally@math.uchicago.edu). Instructors are not permitted to excuse students from the scheduled time of the final exam except in the cases of an Incomplete. The University examination schedule can be found at http://registrar.uchicago.edu/students/final-exams.html.

Homework: Homework should be handed in on or before the given deadline. Late homework will be given zero credit unless an extension is given by me. Extensions and the like will only be given in exceptional circumstances (for example, sudden ill health) and should be requested well before the deadline. Each assigned problem will be graded out of a possible 5 points. A further 20 points will be assigned for the presentation of your work. In order to receive point these points you must do the following. Each question must be answered in \underline{full sentences} (even on the shorter questions). Placing random mathematical symbols around a page is never sufficient to answer a question. An excellent way to check if you work is written up well is to read it back to yourself. You must also hand the questions in in the order in which they are assigned and put \underline{your name} and \underline{your tutors name} on the front sheet, \underline{stapling} the sheets together on the top left-hand corner.

Miscellaneous items: The use of calculators is forbidden, even in doing homework. You are welcome to work on problems together but must write up your solutions independently. Please refer to the University guidelines on plagiarism.

David Rule